A 25.0-kg child in Red Deer pumps herself on a swing when she kicks upward on the downswing, thus changing the distance from the pivot point to her centre of gravity from 2.40 m to 2.28 m.


1.The resonant frequency of her swing before the kick is: ___ Hz. (Give your answer with the correct number of sig digs and do no include units in your answer).


2.The resonant frequency of her swing after the kick is: ___ Hz. (Give your answer with the correct number of sig digs and do no include units in your answer).


1
Expert's answer
2021-07-19T09:52:32-0400

1)


"f_1=\frac{1}{2\pi}\sqrt{\frac{9.8}{2.4}}=0.322\ Hz"

2)


"f_2=\frac{1}{2\pi}\sqrt{\frac{9.8}{2.28}}=0.330\ Hz"